OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y = 2x + 2 and passes through the point (6, 3).

OpenStudy (anonymous):

Perpendicular simply means a line with a negative reciprocal slope. The slope of this equation is "2" or "2/1" The negative reciprocal would be "-1/2" This gives you y=-1/2X + ? Since you have a point the line needs to pass through, you can use the slope to trace back the Y-intercept to find the "?" Answer is y=-1/2X + 6

OpenStudy (anonymous):

thank you !!

OpenStudy (anonymous):

No problem. Unfortunately nobody is solving my problem, as it's too complicated...